OpenStudy (anonymous):

Distance in the Horizon Because of earth's curvature, a person can see a limited distance to the horizon. The higher the location of the person, the farther that person can see. The distance D in miles to the horizon can be estimated by D(h)=1.22 sqrt{h}, where h is the height of the person above ground in feet. a. How high does a person need to be to see 20 miles?

OpenStudy (radar):

It looks like the equation takes care of the conversion of miles and feet, so we can just do it this way:\[20=1.22\sqrt{h}\]\[\sqrt{h}=20/1.22=16.39\] Squaring h=268.745 feet

OpenStudy (radar):

I have seen this equation before, it was used for determining antenna height for "line of sight" communications paths.

OpenStudy (anonymous):

ty so much radar, i am hopeless with this math i swear

OpenStudy (radar):

Keep up the effort and with practice it will be better.

OpenStudy (anonymous):

I hope so, final exams are tuesday, i wish i had a good printout step by step on how to do problems